# FLUID: Training-Free Face De-identification via Latent Identity Substitution

### v1

@@ Abstract
We present FLUID (Face de-identification in the Latent space via Utility-preserving Identity Displacement), a training-free framework that directly substitutes identity in the latent space of pretrained diffusion models. Inspired by substitution mechanisms in chemistry, we reinterpret identity editing as semantic displacement in the latent h-space of a pretrained unconditional diffusion model. Our framework discovers identity-editing directions through optimization guided by novel reagent losses, which supervise for attribute preservation and identity suppression. We further propose both linear and geodesic (tangent-based) editing schemes to effectively navigate the latent manifold. Experimental results on CelebA-HQ and FFHQ demonstrate that FLUID achieves a superior trade-off between identity suppression and attribute preservation, outperforming state-of-the-art de-identification methods in both qualitative and quantitative metrics.
Introduction
@@
The widespread deployment of face recognition (FR) systems has raised urgent concerns about biometric privacy. Facial images are permanent, uniquely identifiable, and easily captured, making them ideal for surveillance but difficult to protect. Although legal safeguards such as the General Data Protection Regulation (gdpr) and India’s Digital Personal Data Protection Act (dpdpa) have emerged, technical solutions remain limited. As a result, de-identification methods are increasingly needed to protect privacy and mitigate risks such as deepfake misuse (westerlund2019emergence) before visual data are shared.
Traditional de-identification techniques (e.g., masking, pixelation, and blurring) (babaguchi2009psychological; ilia2015face) often degrade image quality and fail against modern FR models. Recent generative models, such as generative adversarial networks (GANs) (goodfellow2014generative) and diffusion models (DMs) (ho2020denoising) offer promising alternatives with improved realism (wen2022identitydp; cai2024disguise; shaheryar2025black). However, while these generative methods successfully achieve more realistic identity suppression than traditional ones, they share a critical limitation: most produce visibly altered faces immediately recognizable as de-identified. This occurs because they often fail to preserve well-known non-identity attributes, such as emotion, pose, and gender, resulting in uncanny or artificial-looking outputs. Consequently, viewers may perceive these protected images as unnatural or suspicious, which can further undermine content authenticity and source credibility.
Figure 1: The concept of our approach with the metaphor of substitution reaction. Identity editing within the latent space replaces the original identity (Id) with a de-identified one (Id*), while preserving attributes of a human face image (Hu), such as emotion (Em), gender (Ge), and Pose (Po), strictly than state-of-the-art de-identification methods.
@@
To address these limitations, we turn to latent-space editing, hypothesizing its superior attribute preservation compared to generative reconstruction approaches. Latent editing operates on semantically disentangled representations, allowing independent and gradual manipulation. This enables fine-grained control over the input while reflecting required changes. Recent discoveries in the latent spaces of generative models have enabled such training-free semantic image editing, where inverting input images into the latent space and editing by traversing semantic directions (e.g., emotion, color, pose) yield desired results (harkonen2020ganspace; shen2020interfacegan; kwon2022diffusion). However, current editing methods primarily focus on attribute manipulation while preserving identity, making them unsuitable for de-identification. Moreover, prior editing often relies on inadequate prompts, model fine-tuning, or complex pre-processing pipelines (kim2022diffusionclip; avrahami2022blended; zhu2023boundary; sun2024diffam).
To enable de‑identification with strict attribute preservation, we propose FLUID (Face de‑identification in the Latent space via Utility‑preserving Identity Displacement), a training-free framework that performs identity substitution directly in the latent space of a pretrained unconditional DM. This space, named h-space, of unconditional DMs has been shown to provide a linear, robust, and homogeneous semantic space (kwon2022diffusion). Drawing inspiration from substitution reactions in organic chemistry (rizzo_sn1sn2; ashenhurst_sn1sn2), FLUID metaphorically frames identity removal as a semantic substitution process. In this framework, we introduce a set of reagent losses, each acting as a functional analogue to chemical reagents, to control the outcome of this substitution, specifically targeting identity suppression and attribute preservation. To operationalize this concept, we explore both linear and geodesic (tangent-based) displacement strategies in latent space, enabling us to examine the trade-off between de-identification strength and attribute retention. To the best of our knowledge, we are the first to perform face de-identification in the latent space of a DM. Our contributions are as follows:
  • We introduce a training-free identity substitution framework in the h-space of an unconditional DM.
  • We propose a novel editing scheme inspired by substitution chemistry, optimized via differentiable reagent losses.
@@
  • We demonstrate FLUID’s improved performance over training-based SOTA methods in both realism and utility preservation through extensive experiments.
Figure 2: An illustration of our framework. In the initialization procedure, $x$ is inverted into the latent space of a DM to obtain a starting point $h$, a latent vector in the $h$-space, while three auxiliary models are used to get $F_{x}$, $A_{x}$, and $M$. During optimization, a direction vector $\Delta h$ is iteratively guided by three loss functions to compare $x$ and $\hat{x}$: an identity loss $L_{ID}$, an attribute preservation loss $L_{att}$, and a face mask loss $L_{mask}$. Decoding $\hat{h}$, the combination of $h$ and $\Delta h$, by either linear or tangent edit results in an edited image $\hat{x}$. Note that $\Delta h$ is initialized randomly in the first optimization step.
Related Works
Face Image Editing.
@@
Facial De-identification.
Traditional privacy methods, including blurring and masking (babaguchi2009psychological; ilia2015face), compromise utility and realism. Recent approaches leverage generative models for semantic de-identification, falling into two broad categories: perturbation-based and semantic identity-shifting. Perturbation-based methods, for example, Fawkes (shan2020fawkes), AMT-GAN (hu2022protecting), and DiffAM (sun2024diffam), introduce adversarial noise or stylized makeup to deceive FR models but maintain perceptual identity, limiting protection against human re-identification. Semantic methods like DeepPrivacy2 (hukkelaas2023deepprivacy2), Repaint (lugmayr2022repaint), and IDDiffuse (shaheryar2024iddiffuse) replace facial identity altogether, improving both privacy and transferability. Face-swapping approaches such as DCFace (kim2023dcface) and DiffSwap (zhao2023diffswap) require additional identity inputs, making them vulnerable to misuse in deepfake scenarios (westerlund2019emergence). While most of the methods require training and additional inputs, our method requires a single input and is training-free, achieving effective de-identification.
Preliminaries
Diffusion Models.
@@
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}),$ (1)
@@
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)).$ (2)
@@
$\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right),$ (3)
@@
$x_{t-1}=\sqrt{\alpha_{t-1}}\mathbf{P}_{t}(\epsilon_{\theta}(x_{t}|\tilde{h}_{t}))+\mathbf{D}_{t}(\epsilon_{\theta}(x_{t}))+\sigma_{t}z_{t},$ (4)
where $\tilde{h}_{t}=h_{t}+\Delta h$, $\sigma_{t}$ is the noise scale, and $z_{t}\sim\mathcal{N}(0,\mathbf{I})$. The transformation $\mathbf{P}_{t}$ shifts the $x_{0}$ prediction based on $\tilde{h}_{t}$, while $\mathbf{D}_{t}$ reconstructs the path to $x_{t}$. The key challenge is discovering semantically meaningful directions $\Delta h$ that enable targeted edits without degrading unrelated features.
Proposed Method
This section introduces our method for training-free face de-identification using DMs through latent identity substitution. Following our chemistry-inspired framework, we remove identity information from the original image in the latent space via loss-driven optimization, then substitute a de-identified representation. Our approach operates entirely within the h-space of pretrained unconditional DMs without additional training, consisting of three key components: latent inversion, iterative optimization of identity-editing directions via differentiable losses, and reconstruction through the DM’s decoder. Figure 2 provides an overview of our approach, illustrating the initialization procedure where a source image is inverted into the latent space, followed by the optimization process that uses reagent losses to guide identity substitution while preserving other facial attributes. Unlike direct latent replacement, our method discovers identity-specific directions through optimization rather than assuming prior knowledge of the identity subspace.
Problem Definition
Given a facial image $x\in\mathbb{R}^{H\times W\times 3}$, our goal is to generate a semantically consistent de-identified image $\hat{x}$ such that the identity difference between $x$ and $\hat{x}$ is maximized, i.e.,
@@
where $D_{\text{id}}(\cdot,\cdot)$ is an identity distance function between embeddings. Instead of directly editing in pixel space, we perform identity manipulation in the $h$-space of a pretrained DM. Let $f_{\text{inv}}$ be an inversion function that maps the image to its latent representation:
$z=f_{\text{inv}}(x),\quad z\in\mathbb{R}^{d}.$ (6)
We aim to discover a direction $\Delta h_{\text{id}}\in\mathbb{R}^{d}$ in $h$-space such that applying it to $z$ suppresses identity-related information:
$\hat{z}=z+\lambda\cdot\Delta h_{\text{id}},$ (7)
where $\lambda\in\mathbb{R}$ controls the magnitude of editing. The de-identified image is reconstructed via the DM decoder:
[…]
To facilitate training-free identity editing, we assume that the identity component lies in a locally linear subspace of the h-space. While we do not access this subspace explicitly, we optimize direction vectors guided by differentiable loss signals. Our experimental results support that these assumptions are approximately satisfied, as identity editing can be achieved without degrading other facial attributes. This builds on the empirical properties observed in recent work, such as semantic disentanglement, linear editability, and stable decoder mappings in the bottleneck features of diffusion U-Nets (kwon2022diffusion; jeong2024training). We formalize the notion of latent editability and demonstrate that identity suppression and substitution can be achieved by optimizing a loss in latent space without retraining or fine-tuning the diffusion backbone.
Definition 1.
[…]
Proposition 1.
Let $z\in\mathbb{R}^{d}$ be the latent code of an image $x$, and let $\mathcal{D}$ be a deterministic decoder corresponding to a DM. Define the edited image as:
$\hat{x}=\mathcal{D}(z+\Delta h_{\text{id}}),$ (9)
where $\Delta h_{\text{id}}\in\mathbb{R}^{d}$ is a trainable direction in latent space. Suppose the following conditions hold:
  1. 1.
    Identity attributes are embedded in a low-dimensional affine subspace $\mathcal{H}_{\text{id}}\subset\mathbb{R}^{d}$;
  2. 2.
    The decoder $\mathcal{D}$ is Lipschitz-continuous and differentiable with respect to its input;
  3. 3.
    The latent feature $z$ remains approximately consistent across diffusion denoising timesteps.
Then, there exists an optimal direction $\Delta h_{\text{id}}^{*}\in\mathcal{H}_{\text{id}}$ that minimizes a composite image-space objective:
[…]
with $\hat{x}=\mathcal{D}(z+\Delta h_{\text{id}})$ and scalar weights $\alpha,\beta,\gamma\in\mathbb{R}_{+}$, such that identity information in $\hat{x}$ is suppressed while other facial semantics are preserved.
Sketch of Proof. Because $\mathcal{L}_{\text{total}}$ is differentiable with respect to $\hat{x}$, and $\mathcal{D}$ is differentiable with respect to $\Delta h_{\text{id}}$, the composite gradient is expressed via the chain rule as:
$\nabla_{\Delta h_{\text{id}}}\mathcal{L}_{\text{total}}=\nabla_{\hat{x}}\mathcal{L}_{\text{total}}\cdot\nabla_{\Delta h_{\text{id}}}\mathcal{D}(z+\Delta h_{\text{id}}).$ (11)
Condition 2 ensures that the gradients are well-behaved and bounded. Condition 1 constrains $\Delta h_{\text{id}}$ within a feasible subspace for identity transformation. Finally, condition 3 ensures that the editability holds consistently across intermediate denoising steps, as $z$ remains stable. Therefore, training-free optimization of $\Delta h_{\text{id}}$ in latent space is both theoretically justified and practically realizable using image-level supervision alone.
Latent Navigation: From Euclidean to Geodesic

### v2

@@ Abstract.
Current face de-identification methods that replace identifiable cues in the face region with other sacrifices utilities contributing to realism, such as age and gender. To retrieve the damaged realism, we present FLUID (Face de-identification in the Latent space via Utility-preserving Identity Displacement), a single-input face de-identification framework that directly replaces identity features in the latent space of a pretrained diffusion model without affecting the model’s weights. We reinterpret face de-identification as an image editing task in the latent h-space of a pretrained unconditional diffusion model. Our framework estimates identity-editing directions through optimization guided by loss functions that encourage attribute preservation while suppressing identity signals. We further introduce both linear and geodesic (tangent-based) editing schemes to effectively navigate the latent manifold. Experiments on CelebA-HQ and FFHQ show that FLUID achieves a superior balance between identity suppression and attribute preservation, outperforming existing de-identification approaches in both qualitative and quantitative evaluations.
Face de-identification, face anonymization, utility preservation, diffusion models, image editing, h-space
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Security and privacy Social network security and privacyccs: Computing methodologies Computer visionccs: Computing methodologies Machine learning approaches
1. Introduction
@@
The widespread deployment of face recognition (FR) systems has raised urgent concerns about biometric privacy. Facial images are permanent, uniquely identifiable, and easily captured, making them ideal for surveillance but difficult to protect. In modern digital ecosystems, ranging from public CCTV networks to social media platforms, facial data can be collected, stored, and analyzed at massive scale without an individual’s awareness or consent. Even with legal protections such as the General Data Protection Regulation (gdpr) and India’s Digital Personal Data Protection Act (dpdpa) in place, enforcement is challenging and often lags behind rapid advancements in AI-driven surveillance technologies. As a result, de-identification methods are increasingly needed to protect privacy and mitigate risks such as deepfake misuse (westerlund2019emergence) before visual data are shared. This need becomes especially critical as generative AI enables unprecedented manipulation of facial content, amplifying the potential for misuse if source identities are not adequately protected.
Traditional de-identification techniques (e.g., masking, pixelation, and blurring) (babaguchi2009psychological; ilia2015face) often degrade image quality and fail against modern FR models. Such approaches provide only superficial obfuscation and are easily reversed or bypassed by advanced recognition systems trained on large-scale datasets (radiya-dixit2022data). Prior work has shown that basic obfuscation techniques and data-poisoning cloaks do not offer long-term or robust privacy, as the results are still in danger of identity leakage (oh2016faceless) and adversaries are reversible (nie2022diffusion). Recent generative models, such as generative adversarial networks (GANs) (goodfellow2014generative) and diffusion models (DMs) (ho2020denoising) offer promising alternatives with improved realism (wen2022identitydp; cai2024disguise; shaheryar2025black). These models introduce structured transformations capable of removing identity information while preserving visual coherence, producing images that are in high quality. However, while these generative methods successfully achieve more realistic identity suppression than traditional ones, they still suffer from a fundamental limitation: most produce faces that are immediately recognizable as de-identified. This occurs because they often fail to preserve well-known non-identity attributes, such as gender and age, which leads to outputs that appear artificial or uncanny despite being visually artifact-free. This visible alteration introduces a subtle but important privacy paradox. When an image clearly signals that de-identification has been applied, it undermines privacy rather than strengthening it. First, conspicuous modifications explicitly reveal that sensitive information has been removed, which can attract adversarial scrutiny and motivate targeted attempts to reverse or exploit the anonymization, which would be unlikely if the image appeared authentic. Second, detectably de-identified faces reduce the credibility of visual data in applications where both privacy protection and perceptual authenticity are essential, such as medical case reporting, journalistic documentation, legal evidence, and public research datasets. In real-world deployments, these unnatural distortions do not merely indicate that anonymization has occurred; they can also erode trust in the data itself. An effective face de-identification method should therefore go beyond identity suppression and preserve key semantic cues, producing outputs that remain perceptually indistinguishable from natural, unmodified faces. Only such realism can prevent adversarial suspicion while maintaining the authenticity required for practical, high-stakes data sharing.
@@
To address these limitations, we turn to image editing in the latent space, hypothesizing its superior attribute preservation compared to generative reconstruction approaches. Unlike full image regeneration which often forces a model to re-synthesize identity-irrelevant attributes from scratch, latent editing directly modifies structured internal representations that already encode fine-grained semantic details of the input. Image editing in the latent space operates with semantically disentangled representations, allowing independent and gradual manipulation from the original data. This enables fine-grained control over the input while reflecting required changes. Importantly, such latent-level operations minimize unnecessary modification, making them well-suited for scenarios where utility attributes (e.g., age, gender, emotion, hairstyle) must remain visually consistent. Recent discoveries in the latent spaces of generative models have enabled such training-free semantic image editing, where inverting input images into the latent space and editing by traversing semantic directions (e.g., emotion, color, pose) yield desired results (harkonen2020ganspace; shen2020interfacegan; kwon2022diffusion). These findings collectively suggest that latent spaces possess strong linearity and smooth semantic geometry, allowing meaningful edits without requiring model retraining. However, current face image editing methods primarily focus on attribute manipulation while preserving identity, making them unsuitable for de-identification. Moreover, prior editing often relies on inadequate prompts, model fine-tuning, or complex pre-processing pipelines (kim2022diffusionclip; avrahami2022blended; zhu2023boundary; sun2024diffam). Thus, despite the promise of latent structures, existing pipelines are not directly designed for privacy protection and often lack mechanisms to enforce identity removal while retaining facial utility.
To enable realistic de‑identification with strict attribute preservation, we propose FLUID (Face de‑identification in the Latent space via Utility‑preserving Identity Displacement), which performs identity substitution directly in the latent space of a pretrained unconditional DM without the need to tune model weights. Our core idea is to treat identity removal as a controlled displacement problem in a well-structured semantic manifold, rather than a full generative reconstruction task. This space, named h-space, of unconditional DMs has been shown to provide a linear, robust, and homogeneous semantic space (kwon2022diffusion). In this framework, we introduce a set of loss functions to control the outcome of this process, specifically targeting identity suppression and attribute preservation. These losses collectively enforce a disentangled optimization path: identity information is pushed away, while non-identity features are explicitly constrained to remain stable. As a result, our method aims to generate de-identified faces that remain visually indistinguishable from authentic, unmodified images, achieving what we term ‘covert privacy protection’. By producing outputs that do not visibly signal the application of anonymization, the method avoids attracting adversarial scrutiny and preserves data credibility in scenarios where both privacy protection and perceptual authenticity are essential.
Additionally, we introduce a geodesic (tangent-based) displacement strategy and explore both linear and geodesic edits in latent space, enabling us to examine the trade-off between de-identification strength and attribute retention. This dual-editing view allows us to empirically characterize how different trajectory geometries in latent space influence the final privacy-utility balance. We further assess the realism of results with balanced identity-utility trade-off quantaitatively and qualitatively. To the best of our knowledge, we are the first to perform face de-identification in the latent space of a DM. Our contributions are as follows:
  • We introduce a face de-identification framework in the h-space of an unconditional DM.
@@
  • We demonstrate FLUID’s improved performance over SOTA methods in both realism and utility preservation through extensive experiments.
Figure 1. An illustration of our framework. In the initialization procedure, $x$ is inverted into the latent space of a DM to obtain a starting point $h$, a latent vector in the $h$-space, while three auxiliary models are used to get $F_{x}$, $A_{x}$, and $M$. During optimization, a direction vector $\Delta h$ is iteratively guided by three loss functions to compare $x$ and $\hat{x}$: an identity loss $L_{ID}$, an attribute preservation loss $L_{att}$, and a face mask loss $L_{mask}$. Decoding $\hat{h}$, the combination of $h$ and $\Delta h$, by either linear or tangent edit results in an edited image $\hat{x}$. Note that $\Delta h$ is initialized randomly with a small norm in the first optimization step.
2. Related Works
2.1. Facial De-identification.
Traditional privacy methods, including blurring and masking (babaguchi2009psychological; ilia2015face), provide the best efficiency with simple operations but compromise utility and realism. Recent approaches leverage generative models for semantic de-identification, falling into two broad categories: perturbation-based and semantic identity-shifting. Perturbation-based methods, for example, Fawkes (shan2020fawkes), AMT-GAN (hu2022protecting), and DiffAM (sun2024diffam), introduce adversarial noise or stylized makeup to deceive FR models with subtle changes but maintain perceptual identity, limiting protection against human re-identification. To protect privacy against both humans and FR systems, semantic methods like DeepPrivacy2 (hukkelaas2023deepprivacy2), Repaint (lugmayr2022repaint), and IDDiffuse (shaheryar2024iddiffuse) replace facial identity altogether, improving both privacy and realism. Though face-swapping approaches such as DCFace (kim2023dcface) and DiffSwap (zhao2023diffswap) can perform semantic shift-based de-identification by giving a synthetic identity as target, they require additional identity inputs, making them vulnerable to misuse in deepfake scenarios where a real identity is given to mimic (westerlund2019emergence). Current SOTA semantic shift-based approaches primarily compete and focus on increasing identity dissimilarities between original and de-identified images. While recent advances in generative models have greatly reduced visual artifacts, this aggressive suppression often compromises realism, leading to unintuitive changes in utilities such as age and gender. We argue that the pursuit of extreme identity removal in these approaches sacrifices key utilities, particularly age and gender, thereby undermining the naturalness of de-identified faces that generative methods were originally introduced to improve over traditional techniques. An ideal face de-identification method should instead produce realistic outputs that remain indistinguishable and prevent doubts with further attempts to reverse the de-identification. We tackle this common limitation with a single input without altering the generative model’s weights, while most of the methods require training and additional inputs.
2.2. Face Image Editing.
@@ 2. Related Works
3. Preliminaries
3.1. Diffusion Models.
@@ 3. Preliminaries
(1) $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}),$
@@
(2) $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)).$
@@
(3) $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right),$
@@
(4) $x_{t-1}=\sqrt{\alpha_{t-1}}\mathbf{P}_{t}(\epsilon_{\theta}(x_{t}|\tilde{h}_{t}))+\mathbf{D}_{t}(\epsilon_{\theta}(x_{t}))+\sigma_{t}z_{t},$
where $\tilde{h}_{t}=h_{t}+\Delta h$, $\sigma_{t}$ is the noise scale, and $z_{t}\sim\mathcal{N}(0,\mathbf{I})$. The transformation $\mathbf{P}_{t}$ shifts the $x_{0}$ prediction based on $\tilde{h}_{t}$, while $\mathbf{D}_{t}$ reconstructs the path to $x_{t}$. The key challenge is discovering semantically meaningful directions $\Delta h$ that enable targeted edits without degrading unrelated features. We perform face de-identification in this space than GAN’s latent spaces for two reasons: First, GAN inversion is far from perfect due to its significantly reduced dimensionality from pixel space to latent space (pan2023effective), leading to inaccurate reconstruction. Second, because of the inaccurate GAN inversion, making this inversion accurate requires additional steps, such as fine-tuning the model with a given image (shamshad2023clip2protect) or introducing an additional encoder (maximov2020ciagan), which increases the complexity of the overall image editing process. As diffusion inversion already offers faithful reconstruction and the h-space serves as a semantically meaningful representation, we adopt this space using DDIM inversion (song2020denoising) for effective identity editing.
4. Proposed Method
This section introduces our training-free face de-identification using diffusion models (DMs) through latent identity substitution. We remove identity information from the original image in the latent space via loss-driven optimization, then substitute a de-identified representation. Our approach operates entirely within the h-space of a pretrained unconditional DM, requiring no additional training. It consists of three key components: latent inversion, iterative optimization of identity-editing directions via differentiable losses, and reconstruction through the DM’s decoder. Figure 1 provides an overview of our approach, illustrating the initialization procedure where a source image is inverted into the latent space, followed by the optimization process that uses reagent losses to guide identity substitution while preserving other facial attributes. Unlike direct latent replacement, our method discovers identity-specific edit directions through optimization rather than assuming prior knowledge of an identity subspace.
4.1. Problem Definition
Given a facial image $x\in\mathbb{R}^{H\times W\times 3}$, our objective is to generate a semantically consistent de-identified image $\hat{x}$ such that the identity difference between $x$ and $\hat{x}$ is maximized and the realism perceived by humans is preserved, i.e.,
(5) $),\hat{x}=\operatorname*{arg\,max}_{\!\tilde{x}}(D_{\mathrm{id}}(x,\tilde{x})-D_{\mathrm{human}}(x,\tilde{x}$)),
where $D_{\text{id}}(\cdot,\cdot)$ is an identity distance function between two images and $D_{\text{human}}(\cdot,\cdot)$ is a realism metric determined by human perception. Instead of manipulating the image directly in pixel space (zhu2023boundary; shen2020interfacegan), we perform identity editing in the $h$-space of a pretrained DM. Let $f_{\text{inv}}$ be an inversion function that maps the image to its latent representation:
(6) $z=f_{\text{inv}}(x),\quad z\in\mathbb{R}^{d}.$
Our goal is to discover a direction $\Delta h_{\text{id}}\in\mathbb{R}^{d}$ in $h$-space such that applying this direction to $z$ suppresses identity-related information:
(7) $\hat{z}=z+\lambda\cdot\Delta h_{\text{id}},$
where $\lambda\in\mathbb{R}$ controls the magnitude of editing. Finally the de-identified image is reconstructed using the DM decoder:
@@ 4.1. Problem Definition
To discover identity-specific directions in the h-space, we optimize direction vectors guided by differentiable loss signals. Our experimental results support that face de-identification can be achieved in latent space without degrading other facial attributes. This observation aligns with empirical findings in recent work, such as semantic disentanglement, linear editability, and stable decoder mappings in the bottleneck features of diffusion U-Nets (kwon2022diffusion; jeong2024training). Building on these properties, we formalize the notion of latent editability and show that identity suppression and substitution can be effectively accomplished by optimizing a loss in latent space without retraining or fine-tuning the diffusion backbone.
4.2. Latent Navigation: From Euclidean to Geodesic